Wildhorse Ltd., which has a calendar year end, entered into an equipment lease on June 1, 2020, with GH Financing Limited. The lease term is two years and requires payments of $3,800 at the end of each month beginning September 30. The stated rate of interest in the lease is 6%. As an incentive for entering into the contract, GH has agreed to forgive the first three payments under the lease (June, July, and August).

a.) Calculate the amount that Wildhorse should record for the lease obligation on June 1, 2020. Hint: Calculate the present value of the monthly payments as at September 1, 2020, and then discount this amount to June 1, 2020. (Round answer to 2 decimal places)

b.) What is the amount of the interest accrual that Wildhorse will record on June 30, 2020, for the lease obligation? (Round answer to 2 decimal places)
Answer #1

a. Computation of amount to be recorded for the lease obligation.

\begin{align*}\text{Present Value of Monthly Payments as at September 1, 2020}&=\text{Cash flows}\times \left ( \frac{1-\left ( 1+r \right )^{-n}}{r} \right )\\&=\$3,800\times\left ( \frac{1-\left (1+\frac{0.06}{12} \right )^{-21}}{\frac{0.06}{12}} \right )\\&=\$75,574.32 \end{align*}

Now, discount the above calculated figure and compute the present value as at june 1, 2020

\begin{align*}\text{Present Value as at June 1, 2020}&=\frac{\text{Present value of annuity as at September 1, 2020}}{\left ( 1+r \right )^n}\\&=\frac{\$75,574.32}{\left ( 1+\frac{0.06}{12} \right )^{3}}\\&=\$74,451.94 \end{align*}

b. Computation of Amount of interest accrual that will be recorded on June 30, 2020

\begin{align*}\text{Interest Accrual}&= \text{Lease Obligation}\times \left ( \frac{r}{12} \right )^n\\&=\$74,451.94 \times\left ( \frac{0.06}{12} \right )^1\\&=\$372.26 \end{align*}